OpenStudy (anonymous):

@nikato x-2y=6 3x-6y=18

OpenStudy (anonymous):

are you trying to solve for x & y?

OpenStudy (anonymous):

yes

OpenStudy (shamil98):

divide the second equation by -3, then add them up..

OpenStudy (anonymous):

can i do something with multiplication ?or does it have to be division

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

-3x-6y=18 3x-6y=18 -12y=36 y= -3

OpenStudy (shamil98):

that is wrong. @bere1

OpenStudy (shamil98):

recheck your multiplication.. and yes you can use multiplication, i rather use division as it is faster.

OpenStudy (anonymous):

INFINTE SOLUTIONS ?

OpenStudy (anonymous):

i just saw i did it wrong thank you

OpenStudy (anonymous):

sorry caps its infinite solutions right

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

...

OpenStudy (nikato):

Correct. Infinite solutions. They are the equation of the same line